# DocReward: A Document Reward Model for Structuring and Stylizing Junpeng Liu1 Yuzhong Zhao211footnotemark: 1Bowen Cao1Jiayu Ding3Yilin Jia4Tengchao Lv5 Yupan Huang5Shaohan Huang5Nan Yang5Li Dong5Lei Cui5Tao Ge5Xun Wang5 Huitian Jiao5Sun Mao5FNU Kartik5Si-Qing Chen5Wai Lam1Furu Wei5 1CUHK 2UCAS

### v1

@@ A.3 Source Documents Expansion
To ensure that the reward model learns to assess differences in structure and style rather than content, we applied a rigorous filtering process. Using python-docx, we extracted text from pairs of Microsoft Word DOCX documents and computed their word counts. Only pairs with a word count difference of fewer than 20 words were retained, ensuring comparable content while isolating variation in structure and style. For the constructed training dataset DocPair, both GPT-4o and GPT-5 serve as the base models of agents.
A.4 Prompts
A.5 Details of Extrinsic Evaluation
The Textual Content to Document defined in Figure 4 is adopted as the document agent, with the base model being GPT-5. Three reward models, including random, GPT-5, and DocReward are compared. Once the document agent generates candidates and the reward model selects the top-ranking document from candidates, a highly educated annotator is asked to rank the three documents selected, according to the definitions of professional structure and style defined in section 2. As a result, documents from each reward model are annotated 130 comparison pairs against those of another reward model.
@@ A.5 Details of Extrinsic Evaluation
A.6 Ablation Study of Inputs
Model
Human Preference Accuracy (%)
Synth vs. Synth
Real vs. Synth
Overall
@@ A.6 Ablation Study of Inputs
A.7 More Examples of Case Study
(a) score: 1.92
(b) score: 3.50
(c) score: 5.47
Figure 7: Example 1 of documents with different structures and styles.
(a) score: 2.28
(b) score: 4.26
(c) score: 12.09
Figure 8: Example 2 of documents with different structures and styles.

### v2

@@ A.3 Source Documents Expansion
To ensure that the reward model learns to assess differences in structure and style rather than content, we applied a rigorous filtering process. Using python-docx, we extracted text from pairs of Microsoft Word DOCX documents and computed their word counts. Only synthetic documents with a word count difference of no more than 20 words from the original document and a ROUGE-L score exceeding a threshold are retained, ensuring comparable content while isolating variation in structure and style. For the constructed training dataset DocPair, both GPT-4o and GPT-5 serve as the base models of agents.
A.4 Annotation Protocol and Reliability
Annotation Guidelines.
The detailed guideline for human annotation are presented in Figure 7.
The annotation guidelines consist of general principles that are formulated in an explicit, objective manner. For instance, extremely narrow margins that produce an almost fully saturated page layout are commonly regarded as unprofessional across different cultural and regional contexts.
Figure 7: Detailed guideline for human annotation.
Independence from annotators’ cultural and professional backgrounds.
The annotation was performed by three Ph.D. students with complementary expertise in computer science/math, marketing, and design.
We measured inter-annotator reliability using Cohen’s Kappa; the results are shown in Table 7.
The high agreement indicates that the annotations follow clear, well-defined rules that do not depend on the annotators’ professional training or cultural background, demonstrating the guidelines’ generality and objectivity.
Annotator 1
Annotator 2
Annotator 3
Average
Annotator 1
83.40
80.92
82.15
Annotator 2
83.40
85.90
84.65
Annotator 3
80.92
85.90
83.41
Average
82.15
84.65
83.41
83.40
Table 7: Cohen’s Kappa among annotators.
A.5 Document Types Distribution
The top 30 document types are presented in Figure 8.
Figure 8: Top 30 Document Type Distribution.
A.6 Prompts
A.7 Details of Best-of-N Evaluation
For the Best-of-N experiment, the Textual Content to Document defined in Section 3 is adopted as the document agent, with the base model being GPT-5. Three reward models, including random, GPT-5, and DocReward are compared. Once the document agent generates candidates and the reward model selects the top-ranking document from candidates, a highly educated annotator is asked to rank the three documents selected, according to the definitions of professional structure and style defined in Figure 7. As a result, documents from each reward model are annotated 130 comparison pairs against those of another reward model.
@@ A.7 Details of Best-of-N Evaluation
A.8 Ablation Study of Inputs
In designing the input channels for DocReward, we experimented with two different configurations: a purely visual channel method and a combination method of visual and additional parsing information. The experimental results are summarized in Table 8. Results show that text and bounding box of text span from an additional OCR module are not helpful for the assessment of professional structure and style.
Inputs
Accuracy
@@ A.8 Ablation Study of Inputs
A.9 Case Study
We present a case study on documents with identical textual content but different structures and styles in Figure 9. In case (a), the allocation of whitespace is ineffective, with insufficient space between Last Name and excessive space between First Name, leading to an imbalanced layout. Key fields such as Faculty/Department, Country, and Country Code are not vertically aligned, causing a cluttered and disorganized layout. This poor alignment and inconsistent spacing result in a low score of 1.21 from DocReward. Case (b) adopts a table-like arrangement, but the level-1 heading The teaching staff member is too small and does not stand out from the body text, diminishing its impact. Additionally, the lack of borders around input fields makes it hard to locate items easily, resulting in a moderate score of 2.11. Case (c) provides a clear and well-structured layout, with headings appropriately larger than the body text and better readability, earning the highest rating of 5.34. These results show that DocReward effectively captures document professionalism in structure and style.
(a) score: 1.21
(b) score: 2.11
(c) score: 5.34
Figure 9: Case study: DocReward’s scores reflect structural and stylistic professionalism.